OpenStudy (anonymous):

triangle with a base length of 6 units and a hieght of 7 units

TheSmartOne (thesmartone):

Welcome to OpenStudy! I saw your question in the chat :) Do you still need help? You just asked the question and just closed it too...

OpenStudy (anonymous):

yes i need help please

OpenStudy (anonymous):

i meant i need to find the area of a triangle with a base of 6 units and a height of 7 units

OpenStudy (alexandervonhumboldt2):

area formula s 1/2bh

OpenStudy (alexandervonhumboldt2):

1/2*6*7 solve that

OpenStudy (anonymous):

ty

TheSmartOne (thesmartone):

-_- @AlexandervonHumboldt2 at least let them learn how to solve it instead of giving them all the steps.

OpenStudy (anonymous):

so 21?

OpenStudy (alexandervonhumboldt2):

yesh

TheSmartOne (thesmartone):

So @caine01 The formula to find the area of a triangle is \(\Large A=\frac{1}{2} \times b\times h\)

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @caine01 triangle with a base length of 6 units and a hieght of 7 units \(\color{blue}{\text{End of Quote}}\) And in your question, it tells you that the base is 6 and the height is 7. So we basically substitute the numbers in to the formula!

OpenStudy (anonymous):

yes thank you for the formula

TheSmartOne (thesmartone):

Np. So did you understand how to solve it now? Also if you need any help navigating around the site, then look at this tutorial made for new OpenStudiers! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8